OpenStudy (anonymous):

Can somebody PLEASE help me through this step by step??? I really need to get it finished and get a good grade. PLEASE Show all work .The volume of gas varies directly with temperature and inversely with pressure. Volume is 100m^3 when temperature is 1508 degrees and pressure is 15 lb/cm^2. What is the volume whenthe temperature is 2508 degrees and the pressure is 20 lb/cm^2?

OpenStudy (anonymous):

@tcarroll010 @Agent_Sniffles

OpenStudy (anonymous):

@Callisto @zepdrix @ghazi Please, I'm desperate

OpenStudy (anonymous):

@rebeccaskell94 @him1618 @goformit100 PLEASE HELP

zepdrix (zepdrix):

Were you able to setup the initial equation? `Volume varies directly with Temperature` \[\large V=kT\]The letter \(k\) is a constant of variation. `and inversely with Pressure` \[\large V=kT\left(\frac{1}{P}\right)\]

zepdrix (zepdrix):

Then they give us a bunch of goodies that we can plug into this function in order to solve for \(k\).

zepdrix (zepdrix):

Hmm darn, it looks like we're going to have to be careful with the units also :(

OpenStudy (anonymous):

I know): This is really complicated and I need to have it done within the next 20 minutes. Will you please go step by step and tell me what to do?

zepdrix (zepdrix):

Since we're going to start by solving for \(k\) it might make sense to rewrite our function in terms of V T and P.\[\large V=kT\left(\frac{1}{P}\right)\]Let's multiply both sides by \(P\).\[\large VP=kT\]Divide both sides by \(T\).\[\large k=\frac{VP}{T}\] They told us that when \(V=100m^3\), that \(T=1508^o\) and \(P=15 \frac{lb}{cm^2}\). So let's plug everything in and see what happens.\[\large k=\frac{VP}{T} \qquad \rightarrow \qquad k=\frac{(100m^3)(1508^o)}{(15\frac{lb}{cm^2})}\]

zepdrix (zepdrix):

Hmm so I think we need to be careful with the units. See how we have meters on top, and centimeters on bottom? :OO That's no good. Before we can multiply this all out, we need to fix that.

OpenStudy (anonymous):

Okay, and how do we go about that?

zepdrix (zepdrix):

Err hmm I dunno :\ these units are confusing me .. grr

zepdrix (zepdrix):

Oh I see why I'm getting confused, I plugged those into the wrong spots :) lol

zepdrix (zepdrix):

\[\large k=\frac{VP}{T} \qquad \rightarrow \qquad k=\frac{(100m^3)(15\frac{lb}{cm^2})}{(1508^o)}\]

zepdrix (zepdrix):

So how do we fix that 15? Hmm thinking :\

OpenStudy (anonymous):

Hmmm, I'm not sure/:

zepdrix (zepdrix):

@goformit100 @amistre64 Hmmmmmmmm..

zepdrix (zepdrix):

lol silly gal ^^ try to ask questions earlier XD When you have more time! :3

OpenStudy (anonymous):

I asked it like an hour ago and no one responded! So I had to re-ask it!

OpenStudy (amistre64):

since k is just a multiplier, you can leave it as is and multiply the next set of P/T to it.

OpenStudy (amistre64):

...t/p that is :)

OpenStudy (amistre64):

\[V = k\frac TP\\k=\frac{(100m^3)(15\frac{lb}{cm^2})}{(1508^o)}\] \[V = \frac{(100m^3)(15\frac{lb}{cm^2})}{(1508^o)}\frac TP\] \[V = \frac{(100m^3)(15lb/cm^2)}{(1508^o)}\frac {2508^o}{20lb/cm^2}\]the units cancel out except for m^3 and you should be bale to math whats left